# Evaluating Automated Radiology Report Quality through Fine-Grained Phrasal Grounding of Clinical Findings

### v2

@@ Abstract
Several evaluation metrics have been developed recently to automatically assess the quality of generative AI reports for chest radiographs based only on textual information using lexical, semantic, or clinical named entity recognition methods. In this paper, we develop a new method of report quality evaluation by first extracting fine-grained finding patterns capturing the location, laterality, and severity of a large number of clinical findings. We then performed phrasal grounding to localize their associated anatomical regions on chest radiograph images. The textual and visual measures are then combined to rate the quality of the generated reports. We present results that compare this evaluation metric with other textual metrics on a gold standard dataset derived from the MIMIC collection and show its robustness and sensitivity to factual errors.
@@ 2 Overall approach
The rest of the section describes the above approach in detail.
@@
To extract the anatomical locations from images, we adopted the approach described in [17, 13] that predicts bounding boxes corresponding to a list of 36 anatomical regions cataloged in the chest X-ray lexicon and provided in the ChestImagenome dataset[13]. The bounding boxes for the anatomical regions were detected using a Faster RCNN model trained on labeled bounding boxes in chest X-ray images[13].
Figure 1d shows bounding boxes of the anatomical regions identified in the FFL patterns of the ground truth radiology report in Figure 1c covering anatomical locations of right and left lung, hilar structures, abdomen, and hemidiaphragm. The localization accuracy of the bounding box detector was previously assessed at 0.896 precision and 0.881 recall and was used to reliably generate the ChestImagenome benchmark dataset[13].
@@ 4 Results
We now present results of applying the quality score to assess report quality on a benchmark dataset of chest X-ray images with validated ground truth reports and extracted FFL patterns.
Dataset
For our experiments, we selected the gold dataset of 439 chest x-rays and their ground truth reports from the publicly available clinician validated ChestImagenome[13] collection built from the MIMIC dataset[18]. The dataset also provided FFL patterns covering 60 findings extracted from the finding and impression sections of the ground truth reports to serve for our report quality evaluation. Further, 36 anatomical locations were marked in each of the images and validated by 2 clinicians.
To evaluate report quality, we experimented with open source radiology report generation tools, namely, XrayGPT[14] and RGRG[15], and an internal tool based on GPT-4 being piloted in our hospital. We ran the report generation tool on the benchmark dataset and retained their automated reports. We then extracted the FFL patterns and recorded the bounding boxes of their indicated anatomical locations for the computation of the report quality score.
Evaluation through the proposed measure
@@
Sensitivity of the report quality score

### v3

@@ Abstract
Several evaluation metrics have been developed recently to automatically assess the quality of generative AI reports for chest radiographs based only on textual information using lexical, semantic, or clinical named entity recognition methods. In this paper, we develop a new method of report quality evaluation by first extracting fine-grained finding patterns capturing the location, laterality, and severity of a large number of clinical findings. We then performed phrasal grounding to localize their associated anatomical regions on chest radiograph images. The textual and visual measures are then combined to rate the quality of the generated reports. We present results that compare this evaluation metric with other textual metrics on gold standard datasets.
@@ 2 Overall approach
To extract the anatomical locations from images, we adopted the approach described in [17, 13] that predicts bounding boxes corresponding to a list of 36 anatomical regions cataloged in the chest X-ray lexicon and provided in the ChestImagenome dataset[13]. The bounding boxes for the anatomical regions were detected using a Faster RCNN model trained on labeled bounding boxes in chest X-ray images[13]. Figure 1d shows bounding boxes of the anatomical regions identified in the FFL patterns of the ground truth radiology report in Figure 1c covering anatomical locations of right and left lung, hilar structures, abdomen, and hemidiaphragm. The localization accuracy of the bounding box detector was previously assessed at 0.896 precision and 0.881 recall and was used to reliably generate the ChestImagenome benchmark dataset[13].
@@ 4 Results
We now present results of applying the quality score to assess report quality on a benchmark dataset of chest X-ray images with validated ground truth reports.
Dataset: For our experiments, we selected the gold dataset of 439 chest x-rays and their ground truth reports from the publicly available clinician validated ChestImagenome[13] collection built from the MIMIC dataset[18] and vetted for bias and fairness during their IRB approval. The dataset also provided FFL patterns covering 60 findings extracted from the finding and impression sections of the ground truth reports to serve for our report quality evaluation. Further, 36 anatomical locations were marked in each of the images and validated by 2 clinicians. To evaluate report quality, we experimented with open source radiology report generation tools, namely, XrayGPT[14] and RGRG[15], and an internal tool based on GPT-4 being piloted in our hospital. We ran the report generation tool on the benchmark dataset and retained their automated reports. We then extracted the FFL patterns and recorded the bounding boxes of their indicated anatomical locations for the computation of the report quality score.
Evaluation through the proposed measure:
@@
Sensitivity of the report quality score:
@@
Discussion: Our method exploited the notation of standardized locations for anomalies in chest X-rays. While the metric itself which focuses on both identity and location differences could be used to evaluate phrasal grounding in general scenes, the FFL pattern descriptions may not have such standardized locations. Finally, our method is computationally efficient even for the bipartite matching step due to the small number of regions matched.